if 1+7i÷(2-i).(2-i)=a+bi learn a and b ?
Answer by amarjeeth123(569)

Simplifying and substituting i^2=-1 we get,
1+7i÷4-i^2=a+bi
1+7i÷5=a+bi
Comparing real and imaginary parts we get,
a=1/5
b=7/5